OpenStudy (anonymous):

Please help with this problem... Jerry plans to build a deck that is 189 square feet. What is the Prime factorization of 189? write your answer in exponential form. So i said it was 3 to the 3rd power times 7 and that was right ! but the second part is what is confusing me it says Jerry wants his deck to be more that 6 feet wide. Using your result from above write two possible sets of detentions for Jerrys deck. Thanks so much

OpenStudy (anonymous):

PLEASE HELP ME!!! :)

OpenStudy (anonymous):

someone please help me with this problem because i am real confused and it would really be kind if someone would be nice enough to help me! thanks

OpenStudy (anonymous):

189 = 3 * 63 189 = 3 * 3 * 21 189 = 3 * 3 * 3 * 7 so the prime factorization of 189 is 3 * 3 * 3 * 7 which is 3^3 * 7

OpenStudy (anonymous):

ok so the answer to the second part would be 3 feet * 63 feet and 9 feet * 21 feet?

OpenStudy (anonymous):

Taking 3 * 3 * 3 * 7 the deck could be 9 x 21 ( I found that by 3 * 3 and 3 * 7) 27 x 7 ( found that by 3 * 3 * 3 and 7)

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

but how would i know if that equals 6 feet bc he wants his deck to be 6 feet wide

OpenStudy (anonymous):

?????????????????????